OpenStudy (anonymous):

Estimate the value of the expression by rounding to the given place value. 58 ÷ 5; round 58 to the nearest 10 A. 10 B. 12 C. 14 D. 20

OpenStudy (anonymous):

The answer is B.

OpenStudy (anonymous):

58 / 5 = 11.6. Round that to 12.

OpenStudy (anonymous):

But it says to round 58 to the nearest ten. So 60 / 5 = 12. The answer is still correct.

hero (hero):

\[\frac{58}{5} = \frac{55 + 3}{5} = \frac{55}{5} + \frac{3}{5} = 11 + \frac{3}{5} = 11 + .60 = 11.60 \approx 12\]

OpenStudy (anonymous):

@Hero Yeah but it says to round 58 to the nearest 10. Your answer is still right though.

hero (hero):

Oh well...

hero (hero):

That's more like a hint anyways

OpenStudy (anonymous):

Thank u!

OpenStudy (swag):

See you been getting help this whole time. From one of the best to.

OpenStudy (anonymous):

Lucky for me I have a other question Chosen especially for @SWAG